# Fractional Thouless pumping of solitons: a unique manifestation of bulk-edge correspondence of nonlinear eigenvalue problems

### v1

@@ I Introduction
This work explores bulk-edge correspondence under eigenvalue nonlinearity $\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$, yielding physical observables inaccessible to conventional linear correspondence ($\hat{H}_{\text{ERM}}\Psi=E\Psi$). We implement this through nonlinear topological phase diagrams in an extended Rice-Mele model with next-nearest-neighbor couplings (NNN) $\{t_{a},t_{b}\}$ and nonlinearity strength $g$. First, we present topological phase diagrams for both linear ($H_{\text{ERM}}\Psi=E\Psi$) and nonlinear ($\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$) regimes, showing Chern number $\mathcal{C}$ evolution (Figs. 1(b),(c)). Crucially, we identify fractional Chern numbers in specific parameter regimes (Fig. 1(c)). We then investigate soliton pumping dynamics by tracking center-of-mass evolution through adiabatic cycles by the numerical methods of dynamically evolved soltions and instantaneous solitons with the theoretical framework of nonlinear eigenvalue problem (see Figs. 2-4 ) and anomolous eigenvalue’s nonlinearity problem (see Figs. 5-7) respectively. Both approaches demonstrate how dynamically tuned NNN hoppings and nonlinearity $g$ drive integer-to-fractional Chern phase transitions (Figs. 4(e), 7(e)). Finally, by constructing a 2D momentum-cyclic parameter space, we define topological invariants correlated with nonlinear eigenvalue bifurcations, obtaining a Haldane-model-like phase diagram. Our results predict directly observable fractional nonlinear Thouless pumping in photonic waveguides or cold-atom systems.
@@
The organizational structure of this paper is outlined as follows: In Sec. II, we provide an in-depth introduction to the Hamiltonian and Schrödinger equation of the nonlinear extended RM model. In Sec. III, we investigate the energy spectra and phase diagrams of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions, examining the effects of nonlinearity strength and NNN parameters on these properties. Sec. IV discusses bulk-edge correspondence in the nonlinear eigenvalue regime, elucidating the role of model parameters. Finally, in Sec. V, we summarize the results of our investigation.
@@ II Nonlinear extended Rice-Mele model
The model system studied in this work consists of a one-dimensional interacting bosonic chain with $N$ dimer units. As illustrated in Fig. 1 (a), each unit cell contains two lattice sites (denoted as A and B respectively). As shown in Fig. 1(a), each unit cell comprises two lattice sites (A and B). At the mean-field level, our model system can be well described by the nonlinear extended RM Hamiltonian [74, 75, 76, 77, 78, 79, 80, 56, 81, 82, 78, 83]
@@
In Eq. (2), $\Psi_{n}=[\psi_{n,A},\psi_{n,B}]$ denotes the wave function for the dimer at site $n$ (two atoms per unit cell). The intracell and intercell couplings are $J_{1}=-J-\delta\sin(\omega t)$ and $J_{2}=-J+\delta\sin(\omega t)$ respectively, where $J$ represents the uniform hopping amplitude and $\delta\sin(\omega t)$ describes the time-dependent dimerization order. The staggered sublattice potential $\Delta\cos(\omega t)$ introduces an energy offset between sites, while $g$ quantifies the strength of the focusing Kerr-like nonlinearity [76]. The NNN Hamiltonian $H_{\text{NNN}}$ contains hopping amplitudes $t_{a}$ and $t_{b}$ along sublattices A and B. Crucially, the model reduces to the standard nonlinear Rice-Mele Hamiltonian [76] when $t_{a}=t_{b}=0$. Finally, we note that Hamiltonian (2) provides a direct theoretical framework for recent experiments [56] investigating interaction-induced Thouless pumping in dynamical superlattices or pulsed light propagation in waveguide arrays [76].
@@ II.1 Topological phase diagram for linear and nonlinear eigenvalue problems
Then, we introduce the matrix $P\left(\omega,k\right)$ and $P\left(\omega,k\right)=H_{\text{ERM}}\left(\omega,k\right)-\omega S\left(\omega,k\right)$. The solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (4). In order to gain a more profound understanding of the bulk-edge correspondence model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$, and the equation is transformed into $P\left(\omega,k\right)\varPsi=\lambda\varPsi$. We remark that the auxiliary eigenvalue $\lambda$ does not carry physical significance in the general case, with the notable exception when $\lambda=0$. Hence, the core problem reduces to determining the solution of $P\left(\omega,k\right)\varPsi=\lambda\varPsi$ specifically at $\lambda=0$. At last, the overlap matrix $S$ appearing in Eq. (4), which exhibits a dependence on the nonlinear parameter $\omega$, is constructed as $S=I-M_{S}\sigma_{z}$ with $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ and $\sigma_{z}$ being the $z$-component of Pauli matrix.
@@ II.2 Nonlinear extended Schrödinger equation
In order to investigate the dynamics of soliton, we proceed to obtain equations of motion of nonlinear extended RM Hamiltonian $H_{\text{ERM}}$ by variation of the Hamiltonian $H_{\text{ERM}}$ as $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$ which are described by the following set of nonlinear Schrödinger equations for $j=1,...N-1,N$,
@@
The nonlinear eigenvalue $\mu$ of $\hat{H}_{\text{ERM}}$, defined through $\hat{H}_{\text{ERM}}\Psi=\mu\Psi$, can be derived by substituting $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (5) and (6). We emphasize that Eqs. (5) and (6) represent a form of time-dependent Gross-Pitaevskii equations—equivalently, discrete nonlinear Schrödinger equations. This framework powerfully describes both the dynamics of Bose-Einstein condensates in ultracold quantum gases [56] and pulsed light propagation in waveguide arrays [76].
@@ III Fractional nonlinear Thouless pumping of nonlinear extended RM Hamiltonian
In Sec. II, we presented the nonlinear extended RM Hamiltonian (1) along with the associated discrete nonlinear Schrödinger equations (5) and (6). In Sec. III, we plan to investigate the quantized and fractionally-quantized nonlinear Thouless pumping based on Eqs. (5) and (6). In more details, The objective of subsection III is to present the band structures of the nonlinear extended RM model as well as the pumping behavior of the soliton, while investigating the effects of the parameters $t_{a}$ and $t_{b}$ in the NNN Hamiltonian term (3). In the subsequent subsection (III.2), we adopt the methodology detailed in Ref. [63] to delve into the anomalous nonlinearity of the eigenvalues within the nonlinear extended RM Hamiltonian (1) and its corresponding bulk-edge correspondence.
@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
In this subsection III.1, we are interested into both the quantized nonlinear Thouless pumping and fractional nonlinear Thouless pumping predicted by the topological phase diagram in Figs. 1(b) and (c). This is achieved by numerically solving Eqs. (5) and (6) under a continuous change of the nonlinearity parameter $g$ in Eq. (1) from zero to a finite value. All simulations employ a $100$-unit cell lattice ($200$ sites). The initial soliton state, localized at site $n=0$, is prepared using the iterative self-consistent algorithm from Refs. [73, 76] with trial wavefunction $\Psi_{0}=\text{sech}\left[\lvert x-100\rvert/5\right]$. Here, we use two distinct methodologies to solve Eqs. (5) and (6) numerically. In the first approach entitled the method of dynamically evolved soltions, we use a fourth-order Runge-Kutta algorithm to numerically propagate the initial soliton profile through time evolution governed by Eqs. (5) and (6). This method directly simulates the dynamical evolution of the wavefunction under physical time-dependent driving. In the second approach called by the method of instantaneous solitons, an iterative self-consistent scheme is employed to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation. Under adiabatic conditions ($\omega\rightarrow 0$), both methods yield identical solutions, providing mutual validation of their accuracy.
@@ III.2 Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian
(i) The anomalous eigenvalue’s nonlinearity of nonlinear extended RM Hamiltonian (1) is defined as [73]
@@
Note that Eq. (7) together with $S$ are employed to introduce auxiliary eigenvalues and furthermore the topological edge states of auxiliary eigenstates are topologically inherited as physical edge states, which represents interplay between the topology and nonlinearity of eigenvalues. The overlap matrix $S$ appearing in Eq. (7), which exhibits a dependence on the nonlinear parameter $\omega$, is constructed in the following manner:
@@
with $S_{0}$ being a diagonal matrix $S_{0}=1-\sigma_{z}M_{S}$ and $\sigma_{z}$ the $z$-component of Pauli matrix. Here, $M_{S}\left(\omega\right)$ is defined as $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$.
(ii) Next, we proceed to introduce the matrix $P\left(\omega,t\right)$ as follows:
@@
From Equation (13), the solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (7). In order to gain a more profound understanding of the bulk-edge correspondence model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$,
@@
We remark that the auxiliary eigenvalue $\lambda$ does not carry physical significance in the general case, with the notable exception when $\lambda=0$. Hence, the core problem reduces to determining the solution of Eq. (14) specifically at $\lambda=0$. Before proceeding, it is worth noting two key distinctions in solving Eqs. (4) and (7). First, Eq. (4) describes the eigenvalue problem for the linear Hamiltonian 1 with $g=0$, whereas Eq. (7) corresponds to the nonlinear case with $g\neq 0$. Second, these equations are solved in different representations: Eq. (4) in momentum space and Eq. (7) in coordinate space.
@@ IV Nonlinear bulk-edge correspondence
At last, the NNN parameters $t_{a}$ and $t_{b}$ modify the anomalous eigenvalue solutions of the extended RM model. The most significant modifications occur in the $\omega$-$t$ diagram when $t_{a}=-t_{b}$. Specifically: (i) Nonzero $t_{a}$, $t_{b}$ values generate a new bulk band between existing bands; (ii) When $t_{a}=-t_{b}$, this emergent band vanishes entirely.
@@ V Conclusion and outlook
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Xunzhen Cao, Shujie Cheng, and Biao Wu for stimulating discussions and useful help. We sincerely acknowledge Yucan Yan for his dedicated efforts during the critical inception phase of this project. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12074344).

### v2

@@ I Introduction
This work explores bulk-edge correspondence under eigenvalue nonlinearity $\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$, yielding physical observables inaccessible to conventional linear correspondence ($\hat{H}_{\text{ERM}}\Psi=E\Psi$). We implement this through nonlinear topological phase diagrams in an extended Rice-Mele model with next-nearest-neighbor (NNN) couplings $\{t_{a},t_{b}\}$ and nonlinearity strength $g$. First, we present topological phase diagrams for both linear ($H_{\text{ERM}}\Psi=E\Psi$) and nonlinear ($\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$) regimes, showing Chern number $\mathcal{C}$ evolution (Figs. 1(b),(c)). Crucially, we identify fractional Chern numbers in specific parameter regimes (Fig. 1(c)). We then investigate soliton pumping dynamics by tracking center-of-mass evolution through adiabatic cycles by the numerical methods of dynamically evolved soltions and instantaneous solitons with the theoretical framework of nonlinear eigenvalue problem (see Figs. 2-4 ) and anomalous eigenvalue’s nonlinearity problem (see Figs. 5-7) respectively. Both approaches demonstrate how dynamically tuned NNN hoppings and nonlinearity $g$ drive integer-to-fractional Chern phase transitions (Figs. 4(e), 7(e)). Finally, by constructing a 2D momentum-cyclic parameter space, we define topological invariants correlated with nonlinear eigenvalue bifurcations, obtaining a Haldane-model-like phase diagram. Our results predict directly observable fractional nonlinear Thouless pumping in photonic waveguides or cold-atom systems.
@@
The organizational structure of this paper is outlined as follows: In Sec. II, we provide an in-depth introduction to the Hamiltonian and Schrödinger equation of the nonlinear extended RM model. In Sec. III, we investigate the energy spectra and phase diagrams of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions, examining the effects of nonlinearity strength and NNN parameters on these properties. Sec. IV discusses bulk-edge correspondence in the anomalous eigenvalue’s nonlinearity regime, elucidating the role of model parameters. Finally, in Sec. V, we summarize the results of our investigation.
@@ II Nonlinear extended Rice-Mele model
The model system studied in this work consists of a one-dimensional interacting bosonic chain with $N$ dimer units. As illustrated in Fig. 1 (a), each unit cell contains two lattice sites (denoted as A and B respectively). Within the mean-field approximation, the dynamics of our bosonic dimer chain is effectively captured by a nonlinear extension of the Rice-Mele Hamiltonian, incorporating both NNN couplings, [74, 75, 76, 77, 78, 79, 80, 56, 81, 82, 78, 83]
@@
In above Eqs. (2) and (3), each unit cell consists of two sublattice sites labelled by A and B, and the collective wavefunction $\Psi_{n}=[\psi_{n,A},\psi_{n,B}]$ encapsulates the amplitude on the $n$-th dimer. The intracell and intercell couplings are $J_{1}=-J-\delta\sin(\omega t)$ and $J_{2}=-J+\delta\sin(\omega t)$ respectively, where $J$ sets the uniform hopping amplitude. The staggered sublattice potential $\Delta\sin(\omega t)$ introduces an energy offset between sites, which is a periodically modulates dimerization pattern, while nonlinear interactions are introduced via a Kerr-type term with strength $g$, favoring self-localization [76]. The NNN Hamiltonian $H_{\text{NNN}}$ contains hopping amplitudes $t_{a}$ and $t_{b}$ along sublattices A and B. Crucially, the model reduces to the standard nonlinear Rice-Mele Hamiltonian [76] when $t_{a}=t_{b}=0$. Finally, we note that Hamiltonian (2) provides a direct theoretical framework for recent experiments [56] investigating interaction-induced Thouless pumping in dynamical superlattices or pulsed light propagation in waveguide arrays [76].
@@ II.1 Topological phase diagram for linear and nonlinear eigenvalue problems
We begin by defining the auxiliary matrix $P\left(\omega,k\right)=H_{\text{ERM}}\left(\omega,k\right)-\omega S\left(\omega,k\right)$, whose kernel, $P\left(\omega,\mathbf{k}\right)\varPsi=0$, reproduces the nonlinear eigen-problem stated in Eq. (4). To elucidate the bulk-edge correspondence encoded in this equation, we promote the equality to an eigen-problem $P\left(\omega,k\right)\varPsi=\lambda\varPsi$, with $\lambda\in\mathbb{R}$ serving as a bookkeeping parameter. Physical relevance is restricted to zero-eigenvalue subspace $\lambda=0$, so the central task reduces to solving $P\left(\omega,k\right)\varPsi=\lambda\varPsi$ at $\lambda=0$. The overlap matrix $S$ entering Eq. (4) is constructed as $S=I-M_{S}\sigma_{z}$, where $\sigma_{z}$ donates the Pauli z-matrix and $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ encodes the nonlinear $\omega$-dependence.
@@ II.2 Nonlinear extended Schrödinger equation
To study soliton dynamics, we derive the equations of motion for the nonlinear extended RM Hamiltonian $H_{\text{ERM}}$ by applying the variational principle $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$, yielding the following discrete nonlinear Schrödinger equations for sites $j=1,...N-1,N$,
@@
Inserting the ansatz $\Psi(t)=\Psi e^{i\mu t/\hbar}$ into Eqs. (5) and (6) yields a sationary nonlinear eigenvalue problem $\hat{H}_{\text{ERM}}\Psi=\mu\Psi$. Eqs. (5) and (6) constitute a set of discrete nonlinear Schrödinger equations, formally equivalent to the Gross-Pitaevskii framework in lattice systems. This formalism effectively models both ultracold-atom systems [56] and nonlinear photonic lattices under periodic driving [76].
@@ III Fractional nonlinear Thouless pumping of nonlinear extended RM Hamiltonian
In Sec. II, we presented the nonlinear extended RM Hamiltonian (1) along with the associated discrete nonlinear Schrödinger equations (5) and (6). In Sec. III, we plan to investigate the quantized and fractionally-quantized nonlinear Thouless pumping based on Eqs. (5) and (6). In more details, The objective of subsection III is to present the band structures of the nonlinear extended RM model as well as the pumping behavior of the soliton, while investigating the effects of the parameters $t_{a}$ and $t_{b}$ in the NNN Hamiltonian term (3). In the subsequent subsection (III.2), we adopt the methodology detailed in Ref. [63] to delve into the anomalous eigenvalue’s nonlinearity within the nonlinear extended RM Hamiltonian (1) and its corresponding bulk-edge correspondence.
@@ III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
In this subsection III.1, we are interested into both the quantized nonlinear Thouless pumping and fractional nonlinear Thouless pumping predicted by the topological phase diagram in Figs. 1(b) and (c). This is achieved by numerically solving Eqs. (5) and (6) under a continuous change of the nonlinearity parameter $g$ in Eq. (1) from zero to a finite value. All simulations employ a $100$-unit cell lattice ($200$ sites). The initial soliton state, localized at site $n=0$, is prepared using the iterative self-consistent algorithm from Refs. [73, 76] with trial wavefunction $\Psi_{0}=\text{sech}\left[\lvert x-100\rvert/5\right]$. Here, we use two distinct methodologies to solve Eqs. (5) and (6) numerically. In the first approach entitled the method of dynamically evolved soltions, we use a fourth-order Runge-Kutta algorithm to numerically propagate the initial soliton profile through time evolution governed by Eqs. (5) and (6). This method directly simulates the dynamical evolution of the wavefunction under physical time-dependent driving. In the second approach called by the method of instantaneous solitons, an iterative self-consistent scheme is employed to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation. In the adiabatic regime ($\omega\rightarrow 0$), the two approaches converge to the same soliton trajectory, confirming numerical consistency.
@@ III.2 Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian
(i) The nonlinear extended RM Hamiltonian (1) is characterized by an anomalous form of eigenvalue nonlinearity, expressed as [73]
@@
Noted that Eq. (7) and the accompanying matrix S serve to embed auxiliary eigenvalues into the problem; As a consequence, the topological edge modes of these auxiliary eigenstates are inherited by the physical system, thereby highlighting a subtle interplay between eigenvalue nonlinearity and topology. The overlap matrix S in Eq. (7), whose structure varies with the nonlinear parameter $\omega$, is assembled as follows:
@@
with the diagonal block $S_{0}$ takes the form $S_{0}=1-\sigma_{z}M_{S}$ and $\sigma_{z}$ the $z$-component of Pauli matrix. Here, the scalar function $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ governs its dependence on the nonlinear parameter $\omega$.
(ii) We then construct the matrix $P\left(\omega,t\right)$ as follows:
@@
Equation (13) shows that any solution of $P\left(\omega,\mathbf{k}\right)\varPsi=0$ also satisfies the nonlinear relation in Eq. (7). To explore the bulk-edge correspondence implied by this relation more deeply, we append an auxiliary, real-valued eigenvalue $\lambda$, so that
@@
Although the auxiliary eigenvalue $\lambda$ lacks direct physical meaning in general, the special case $\lambda=0$ is singled out. Thus, the central task becomes solving Eq. (14) precisely at this point. Before proceeding, it is worth noting two key distinctions in solving Eqs. (4) and (7). First, Eq. (4) describes the eigenvalue problem for the linear Hamiltonian 1 with $g=0$, whereas Eq. (7) corresponds to the nonlinear case with $g\neq 0$. Second, these equations are solved in different representations: Eq. (4) in momentum space and Eq. (7) in coordinate space.
@@ IV Nonlinear bulk-edge correspondence
At last, the NNN parameters $t_{a}$ and $t_{b}$ modify the anomalous eigenvalue solutions of the extended nonlinear RM model. The most significant modifications occur in the $\omega$-$t$ diagram when $t_{a}=-t_{b}$. Specifically: (i) Nonzero $t_{a}$, $t_{b}$ values generate a new bulk band between existing bands; (ii) When $t_{a}=-t_{b}$, this emergent band vanishes entirely.
@@ V Conclusion and outlook
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Xuzhen Cao, Shujie Cheng, and Biao Wu for stimulating discussions and useful help. We sincerely acknowledge Yucan Yan for his dedicated efforts during the critical inception phase of this project. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12074344).